OpenStudy (anonymous):

What is the value of y in the parallelogram below? If necessary, round your answer to two decimal places

OpenStudy (anonymous):



OpenStudy (anonymous):

well the 4 corners in a 4 sided polygon add up to 360. 1 of them is 360/4=90. so the diagonals should equal 180 so 65+(2x+3)=180. solve for x

OpenStudy (anonymous):

The 4 corners of a parallelogram equal 360 degrees. Opposite corners are equal, and J+K=180

OpenStudy (anonymous):

how do i solve for x?

OpenStudy (anonymous):

According to your question, you aren't solving for x. You're solving for y

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

looks wrong luis